Staff:

LBTO: Guido, Xianju, Brandon and Juan Carlos

Summary.

  • Running the system at1700Hz to check close loop performances. Not issues to report for this setup
  • Frequency 1350Hz showed in close loop with gain0.110modi the "fake" pie shape.
  • No z0 data acquisition done.
  • Further analysis of the "fake" pie and the 127Hz vibration proposed.
  • Take time to run a pupil wobble montioring. Pupils in range
  • During the close loop for pupil wobble the system wqs stable and not issues reported.

Details.

8:20am Power on WFS.

Argos Cal unit reported an error from the OSS.

Looking at the driver is was in error state and the requirements were not fullfille

Set flat.

table_LUCIFER_ACE-AO.txt.20221012_testz0

table_LUCIFER_ACE-AO.txt.20221012_testz0

TN:20221013_161551

gain: 1.40,0.35,0.43

_162032,_162125

gopt.

0.69,0.17,0.21

162324,_162342,

loop at 1350.

fake pi shape.

skipping frame.

172719

20K frames

_172838,

10:36am

adsec rest/flat.

11:09 AM

Pupil wobble check.

Bin1

rot@340

CL

x->26.67

y->27.05

start 11:22

11:40 am

Bin4

preset 16.7

rot 360

CL:

X->40

y->51

start 11:43am

forward: 1200

bin3

shutter 48

mag 16.5

12:16PM

bin2

  • shutter 55
  • mag 14
  • 12:38PM
  • rotA@341 cl: y->50, x->21

-- JuanCarlosGuerra - 13 Oct 2022
Topic revision: r2 - 17 Oct 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback